Circular No. 6943

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA OPHIUCHI 1998 A. V. Filippenko, D. C. Leonard, and M. Modjaz, University of Cali-

fornia at Berkeley; and R. G. Eastman, Lawrence Livermore National Laboratory, report that CCD spectra (range 320–800 nm, resolution 0.6 nm) obtained on June 18 UT with the Shane 3-m reflector at Lick Observatory reveal that the object reported by Nakano and Takamizawa on IAUC 6941 is indeed a nova. Strong emission lines of H and He, as well as the Bowen feature near 465 nm, are superposed on a featureless continuum. The strongest lines (especially Hα) exhibit a broad, box-shaped base (FWZI nearly 12 000 km/s) and a narrow core (FWZI ∼ 2000 km/s).

K. Takamizawa reports that the limiting magnitude of his patrol films (cf. IAUC 6941) is ∼ 15. D. Hanžl, Brno, Czech Republic, reports the following position obtained with a 0.40-m reflector (+ CCD) on June 17: α = 17^{h}31^{m}59^{s}.79, δ = −19^{o}13′56′′.0 (equinox 2000.0). R. A. Kowalski, Zephyrhills, FL, reports the following position end figures and magnitude obtained with a 0.18-m reflector (+ unfiltered CCD) on June 18.244 UT: 59^{s}.78, 55′′.9, 10.4. CCD photometry by Hanžl, based on comparison star GSC 6243.435 (Tycho magnitudes V = 10.71 ± 0.09, B − V = +0.892 ± 0.16): June 17.971 UT, V = 10.32 ± 0.02, B − V = +0.68 ± 0.05.

SUPERNOVA 1998cm IN ANONYMOUS GALAXY L. Germany reports, on behalf of the Mount Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 6639), the discovery of an apparent supernova (V = 18.6) on V and R CCD images taken on June 10 (limiting mag V = 19.0) by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1998cm is located near Abell 1780 at α = 13^{h}44^{m}12^{s}.97, δ = +3^{o}00′40′′.3 (equinox 2000.0), which is 0′′.2 east and 0′′.4 north from the center of its host galaxy. A nearby star has position end figures 15^{s}.27, 5′′.7. SN 1998cm was not visible on similar images taken on May 18 (limiting mag V = 19.0) but was confirmed by S. Chan with the same telescope on June 14 (limiting mag V = 19.0).

A. V. Filippenko, D. C. Leonard, and M. Modjaz, University of California at Berkeley; and R. G. Eastman, Lawrence Livermore National Laboratory, report that CCD spectra (range 320–1000 nm) obtained on June 18 UT with the Shane 3-m reflector at Lick Observatory reveal that the object is a supernova of type Ia, probably within a week or two of maximum brightness. The approximate redshift is 0.08, based on narrow emission lines in the host galaxy.

1998 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green